OpenStudy (anonymous):

How do I do this? Please help. Find tan θ if sec θ =sqrt(37)/6 and sin θ < 0.

OpenStudy (anonymous):

@Darion we will use the formula sec^2(theta) - tan^2(theta)= 1 sec^2(theta) - 1 = tan^2(theta) now u need to put the values given in the ques

OpenStudy (anonymous):

so I replace theta with the value from the question?

OpenStudy (anonymous):

yah ,in ques, value of sec(theta) is given u need to put the value here

OpenStudy (anonymous):

@darion any confusion?

OpenStudy (anonymous):

So it would be sec^2(sqrt(37)/6) - 1 = tan^2(theta)?

OpenStudy (anonymous):

nope the formula is sqr{sec(theta)} - 1 = sqr{tan(theta)} u r given the value of sec(theta) = sqrt(37)/6

OpenStudy (anonymous):

after putting the value we get sqr(sqrt(37)/6) - 1 = sqr{tan(theta)} 37/36-1 = sqr{tan(theta)} 1/36 = sqr{tan(theta)} 1/6 = tan(theta)

OpenStudy (anonymous):

@darion right??????

OpenStudy (anonymous):

Oh. I guess that makes sense. Would it be -1/6?

OpenStudy (anonymous):

yup tan(theta) =±1/6

OpenStudy (anonymous):

in this ques tan(theta) would be equal to -1/6 bcoz it is given sin(theta) less than 0 from this it is concluded that range of sine function is [-1,0) as a result of this we would choose the value which should be less than 0 bcoz range of tangent function is R (real numbers)